### v1

# An Explicit Entire Function of Order One with All Zeros on a Line and Bounded in a Half-Plane

Abstract.
We construct a single explicit entire function $\Xi_{c}(s)$ of order 1, with all zeros provably on $\operatorname{Re}(s)=1/2$, satisfying a functional equation $\Xi_{c}(s)=\Xi_{c}(1-s)$, whose normalized form $Z_{c}(s)=\Xi_{c}(s)/[\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)]$ is uniformly bounded for $\operatorname{Re}(s)>1+\delta$ (any $\delta>0$) yet satisfies $\sup_{t}|Z_{c}(1+it)|=+\infty$. The function thus satisfies an analogue of the Riemann Hypothesis together with the sharp bounded/unbounded transition at $\sigma=1$ characteristic of $\zeta$. The construction uses Hadamard products with controlled zero perturbations; the transition is proved unconditionally via a dyadic large-sieve argument and a Phragmén–Lindelöf transfer.
1. Introduction
The Riemann zeta function exhibits a sharp transition at $\sigma=1$:
  • For $\sigma>1$: $|\zeta(\sigma+it)|\leq\zeta(\sigma)<\infty$.
  • On $\sigma=1$: $\limsup_{t\to\infty}|\zeta(1+it)|=\infty$, with $|\zeta(1+it)|=\Omega(\log\log t)$ [14, Thm. 8.5].
The Selberg class [12, 9] axiomatizes $L$-functions via five properties: Dirichlet series, analytic continuation, functional equation, Ramanujan bound, and Euler product. These axioms are top-down: they prescribe what $L$-functions should look like, based on the structure of $\zeta$ and its generalizations. But they do not reveal which properties are truly necessary for the Riemann Hypothesis or the $\sigma=1$ transition. The Euler product is widely regarded as essential [7, 13]: without it, the Davenport–Heilbronn function [4] has zeros off the critical line, and Bombieri–Garrett’s pseudo-Laplacians [2] fail to capture all zeros spectrally. The de Branges space framework [6] provides a natural setting for entire functions with real zeros, but verifying the required positivity conditions is equivalent to RH itself. For general background on $L$-functions and the analytic tools used here, see [8].
A natural question is whether one can construct a function that provably satisfies an analogue of RH together with the $\sigma=1$ transition. For $\zeta$ itself, the location of zeros is the central open problem; but for a suitably designed function, one can hope to build in the zeros-on-the-line property while retaining enough structure for the $\sigma=1$ transition to hold.
This paper provides such a construction: an explicit entire function $\Xi_{c}(s)$ of order $1$ with all zeros provably on $\operatorname{Re}(s)=1/2$, satisfying $\Xi_{c}(s)=\Xi_{c}(1-s)$, whose normalized form is uniformly bounded for $\sigma>1$ yet unbounded on $\sigma=1$—the same sharp dichotomy exhibited by $\zeta$. The associated ratio $Z_{c}/Z_{0}$ is a Dirichlet series with generalized Euler product, providing a concrete “RH-satisfying” $L$-function–like object whose analytic properties can be studied unconditionally.
Theorem 1.1 (Main Theorem).
For any $c>0$, define $\Xi_{c}(s)$ as in §2, with $Z_{c}(s)=\Xi_{c}(s)/[\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)]$. Then:
  1. (a)
    All zeros of $\Xi_{c}$ lie on $\operatorname{Re}(s)=1/2$.
  2. (b)
    $\Xi_{c}(s)=\Xi_{c}(1-s)$.
  3. (c)
    The zero-counting function satisfies $N_{c}(T)=\frac{T}{2\pi}\log\frac{T}{2\pi}-\frac{T}{2\pi}+O(\log T)$.
  4. (d)
    For each $\delta>0$, $\sup_{t}|Z_{c}(1+\delta+it)|\leq\exp(c\,\zeta_{\alpha}(1+\delta)+O(1))$.
  5. (e)
    $\sup_{t}|Z_{c}(1+it)|=+\infty$.
Parts (a)–(c) hold by construction; (d) via the spectral identity (Proposition 3.1) with linearization error controlled by Lemma 1.2; and (e) via a Phragmén–Lindelöf argument (§5). Moreover, the ratio $Z_{c}/Z_{0}$ is itself a Dirichlet series with generalized Euler product, bounded for $\sigma>1$ and unbounded on $\sigma=1$, with all zeros provably on $\sigma=1/2$.
Lemma 1.2 (Linearization convergence).
(1) $\sum_{n=1}^{\infty}\frac{E(\gamma_{n}^{(0)})^{2}}{n^{2}}<\infty,$
where $E$ is the full (infinite) series (3) evaluated at the regular zero positions.
Proof.
Decompose $\mathbb{N}$ into dyadic blocks $I_{j}=[2^{j},2^{j+1})$ for $j\geq 1$. It suffices to show $\sum_{j}S_{j}<\infty$ where $S_{j}=\sum_{n\in I_{j}}E(\gamma_{n}^{(0)})^{2}/n^{2}$. For each $j$, set $K_{j}=\lceil 2^{1.1\,j}\rceil$ and split $E=E_{K_{j}}+R_{K_{j}}$, where $E_{K_{j}}$ is the first $K_{j}$ terms. By $(a+b)^{2}\leq 2a^{2}+2b^{2}$: $S_{j}\leq 2(\text{HEAD}_{j}+\text{TAIL}_{j})$.
Head (large sieve). $E_{K_{j}}$ is a trigonometric polynomial with $K_{j}$ terms and frequency range $\Lambda_{j}=\omega_{K_{j}}-\omega_{1}=O(j)$. The zeros $\gamma_{n}^{(0)}$ have spacing $\delta\geq c_{1}/j$ for $n\leq 2^{j+1}$. By the large sieve inequality for trigonometric polynomials at well-spaced points (cf. [10, Thm. 1], [11, §7]):
$\sum_{n\leq 2^{j+1}}\!|E_{K_{j}}(\gamma_{n}^{(0)})|^{2}\leq(\pi/\delta+\Lambda_{j})\,\tfrac{1}{\pi^{2}}\sum_{k\leq K_{j}}\alpha_{k}^{-1}\leq C_{1}\,j\log j,$
since $\sum_{k\leq K_{j}}\alpha_{k}^{-1}\sim\log\log p_{K_{j}}\sim\log j$ (Mertens). Therefore $\text{HEAD}_{j}\leq C_{1}\,j\log j\,/\,4^{j}$, and $\sum_{j}j\log j/4^{j}<\infty$.
Tail (Abel summation). From the proof of Lemma 2.1, the tail $D_{K}=\sum_{k>K}(\omega_{k+1}-\omega_{k})/\alpha_{k}^{1/2}$ is a left Riemann sum for the decreasing function $e^{-\omega/2}$. By Bertrand’s postulate (mesh width $\leq\log 2$): $D_{K}\leq\sqrt{2}\int_{\omega_{K}}^{\infty}e^{-\omega/2}\,d\omega=2\sqrt{2}\,\alpha_{K}^{-1/2}\leq 2\sqrt{2}/\sqrt{p_{K}}$. Hence $|R_{K_{j}}(\gamma_{n}^{(0)})|\leq(C_{\theta}/\pi)(\alpha_{K_{j}+1}^{-1/2}+\gamma_{n}D_{K_{j}})\leq C_{2}\,2^{j(1-\varepsilon)/2}/j^{3/2}$ (with $\varepsilon=0.1$, using $\gamma_{n}\leq C_{3}\,2^{j}/j$ and $\sqrt{p_{K_{j}}}\geq C_{4}\,2^{(1+\varepsilon)j/2}\sqrt{j}$). Therefore $\text{TAIL}_{j}\leq C_{5}/(j^{3}\cdot 2^{\varepsilon j})$, and $\sum_{j}1/(j^{3}\cdot 2^{\varepsilon j})<\infty$.
Both contributions decay exponentially, so $\sum S_{j}<\infty$. ∎
1.1. Method
Given a sequence of positive reals $0<t_{1},t_{2},\ldots$ with $\sum t_{n}^{-2}<\infty$, define the Hadamard product
(2) $\Xi(s)=\prod_{n=1}^{\infty}\left(1+\frac{(s-\tfrac{1}{2})^{2}}{t_{n}^{2}}\right).$
This has zeros at $s=1/2\pm it_{n}$ and satisfies $\Xi(s)=\Xi(1-s)$ (cf. [5, §1.10], [14, §2.12]). We choose the zero positions $\{t_{n}\}$ so that the counting function $N_{c}(T)=\#\{n:t_{n}\leq T\}$ approximates $N_{\mathrm{main}}(T)+cE(T)$, where $E$ is a trigonometric series whose Fourier amplitudes decay like the prime reciprocals.
The growth of $|Z|$ on vertical lines is controlled by the Poisson extension of $E$: bounded for width $w>1/2$ (i.e., $\sigma>1$) and unbounded at width $w=1/2$ (i.e., $\sigma=1$), from a marginal convergence/divergence condition on the damped Fourier amplitudes.
2. The Explicit Construction
2.1. Parameters
Fix an irrational $\theta$ (e.g., $\theta=e^{\sqrt{2}}$), an amplitude $c>0$, and the primes $p_{1}<p_{2}<\cdots$. Define:
(3) $\alpha_{k}=p_{k},\quad\omega_{k}=\log p_{k},\quad\phi_{k}=k\theta,\quad E(T)=-\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(\omega_{k}T+\phi_{k})}{p_{k}^{1/2}}.$
Lemma 2.1 (Convergence).
The series $E(T)$ converges for all $T>0$.
Proof.
Write $E(T)=-(1/\pi)\operatorname{Im}\bigl[\sum_{k}d_{k}\,c_{k}\bigr]$ where $c_{k}=e^{ik\theta}$ and $d_{k}=e^{i\omega_{k}T}/\alpha_{k}^{1/2}$. Apply summation by parts (Abel’s formula): $\sum_{k=1}^{K}c_{k}d_{k}=G_{K}d_{K}-\sum_{k=1}^{K-1}G_{k}(d_{k+1}-d_{k})$, where $G_{k}=\sum_{j=1}^{k}e^{ij\theta}$ satisfies $|G_{k}|\leq 2/|1-e^{i\theta}|=:C_{\theta}$ (since $\theta$ is irrational). The boundary term satisfies $|G_{K}d_{K}|\leq C_{\theta}\,\alpha_{K}^{-1/2}\to 0$. For the variation sum: $|d_{k+1}-d_{k}|\leq|\omega_{k+1}-\omega_{k}|\,T/\alpha_{k}^{1/2}+(\alpha_{k}^{-1/2}-\alpha_{k+1}^{-1/2})$. The amplitude differences telescope: $\sum(\alpha_{k}^{-1/2}-\alpha_{k+1}^{-1/2})=\alpha_{1}^{-1/2}$. The phase differences satisfy $\sum_{k}(\omega_{k+1}-\omega_{k})/\alpha_{k}^{1/2}=\sum_{k}\log(\alpha_{k+1}/\alpha_{k})/\alpha_{k}^{1/2}<\infty$ (this sum is a Riemann sum for $\int_{\omega_{1}}^{\infty}e^{-\omega/2}\,d\omega<\infty$, with bounded mesh by Bertrand’s postulate). Hence $\sum|G_{k}(d_{k+1}-d_{k})|$ converges absolutely, and $G_{K}d_{K}\to 0$, so $\sum c_{k}d_{k}$ converges. ∎
2.2. The perturbed zeros
Define the smooth counting function
(4) $N_{\mathrm{main}}(T)=\frac{T}{2\pi}\log\frac{T}{2\pi}-\frac{T}{2\pi}+\frac{7}{8},\qquad N_{\mathrm{main}}^{\prime}(T)=\frac{1}{2\pi}\log\frac{T}{2\pi}.$
Let $\{\gamma_{n}^{(0)}\}$ be the “regular” sequence defined by $N_{\mathrm{main}}(\gamma_{n}^{(0)})=n$, with Hadamard product $\Xi_{0}(s)$. Define the “perturbed” sequence $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$ with
(5) $\delta_{n}=-\frac{cE(\gamma_{n}^{(0)})}{N_{\mathrm{main}}^{\prime}(\gamma_{n}^{(0)})},$
and let $\Xi_{c}(s)$ be the Hadamard product (2) with $t_{n}=\gamma_{n}$.
Lemma 2.2 (Well-definedness of the Hadamard product).
For any $c>0$, the sequence $\{\gamma_{n}\}$ satisfies $\gamma_{n}>0$ for all $n$ (taking $c$ small enough for finitely many small $n$) and $\sum\gamma_{n}^{-2}<\infty$, so $\Xi_{c}(s)$ is an entire function of order $1$.
Proof.
Since $E(T)$ converges for each $T$ (Lemma 2.1), $\delta_{n}$ is finite for each $n$. For $n$ large, $|\delta_{n}|=O(c|E(\gamma_{n}^{(0)})|/\log\gamma_{n}^{(0)})$. By the Abel-summation bound of Lemma 2.1, $|E(T)|=O(T)$ for each $T$ (the bound grows at most linearly since the total variation of $e^{i\omega_{k}T}$ over $k\leq K$ is $O(T\log\alpha_{K})$, controlled by the monotone decay of $\alpha_{k}^{-1/2}$). Thus $|\delta_{n}|=O(c\gamma_{n}^{(0)}/\log\gamma_{n}^{(0)})<\gamma_{n}^{(0)}/2$ for $n$ sufficiently large, giving $\gamma_{n}>\gamma_{n}^{(0)}/2>0$. Since $\gamma_{n}^{(0)}\sim 2\pi n/\log n$, we have $\sum\gamma_{n}^{-2}\leq 4\sum(\gamma_{n}^{(0)})^{-2}<\infty$. The exponent of convergence equals $1$. ∎
Remark 2.3.
The zeros $\gamma_{n}$ need not be in increasing order: crossings ($\gamma_{n+1}<\gamma_{n}$) may occur where $|E^{\prime}|$ is large. For a finite truncation $E_{K}$, global monotonicity of $F_{K}(T)=N_{\mathrm{main}}(T)+cE_{K}(T)$ holds for $c<\log(T_{0}/(2\pi))/(2\pi C_{K})$ with $C_{K}=\frac{1}{\pi}\sum_{k=1}^{K}\omega_{k}\alpha_{k}^{-1/2}$. For the infinite series, occasional crossings produce zeros of multiplicity >1>1 without affecting the Hadamard product convergence or the spectral identity (which depends only on the linearization at $\gamma_{n}^{(0)}$, not on the ordering of $\gamma_{n}$).
3. The Spectral Identity
The following proposition relates the Hadamard product ratio $\Xi_{c}/\Xi_{0}$ to the Fourier content of $E$. All quantities ($\alpha_{k},\omega_{k},\phi_{k},E$) are as defined in §2.
Proposition 3.1 (Spectral Identity).
For $\sigma\geq 1$ and $|t|\geq 2$:
(6) $\log\frac{|\Xi_{c}(\sigma+it)|}{|\Xi_{0}(\sigma+it)|}=-c\sum_{k}\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c).$
Proof.
Step 1 (Linearization). By (5), $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$. Each Hadamard factor satisfies (Taylor-expanding in $\delta_{n}$):
$\log\frac{1+(s-\tfrac{1}{2})^{2}/\gamma_{n}^{2}}{1+(s-\tfrac{1}{2})^{2}/(\gamma_{n}^{(0)})^{2}}=\frac{-2w^{2}\delta_{n}}{\gamma_{n}^{(0)}((\gamma_{n}^{(0)})^{2}+w^{2})}+O(\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2})$
where w=s1/2w=s-1/2.
Step 2 (Finite-truncation limit). For each $K$, let $\Xi_{c,K}$ be the Hadamard product with zeros defined using the truncation $E_{K}$ (first $K$ terms of $E$). The perturbations $\delta_{n,K}=-cE_{K}(\gamma_{n}^{(0)})/N_{\mathrm{main}}^{\prime}(\gamma_{n}^{(0)})$ satisfy $|\delta_{n,K}|\leq 2\pi cM_{K}/\log\gamma_{n}^{(0)}$ where $M_{K}=\frac{1}{\pi}\sum_{k=1}^{K}\alpha_{k}^{-1/2}$ is a constant. The linearization error for the $K$-truncation is then
(7) $\log\frac{\Xi_{c,K}(s)}{\Xi_{0}(s)}=-c\int_{\gamma_{1}^{(0)}}^{\infty}\frac{2w^{2}\,E_{K}(u)}{u(u^{2}+w^{2})}\,du+O(c^{2}),$
where $O(c^{2})$ accounts for the quadratic remainder $\sum\delta_{n,K}^{2}/(\gamma_{n}^{(0)})^{2}$, which converges since $|\delta_{n,K}|=O(cM_{K}/\log\gamma_{n}^{(0)})$ and $\sum 1/(\gamma_{n}^{(0)}\log\gamma_{n}^{(0)})^{2}<\infty$. This error is $O(c^{2}M_{K}^{2})$, independent of $t$.
Since $E_{K}\to E$ pointwise and $\gamma_{n,K}\to\gamma_{n}$ for each $n$, the Hadamard products converge: $\Xi_{c,K}(s)\to\Xi_{c}(s)$ for each $s$, by dominated convergence of the product factors. Taking $K\to\infty$ in (7):
Step 3 (Contour integration). For a single component $E_{k}(u)=-\frac{1}{\pi}\alpha_{k}^{-1/2}\sin(\omega_{k}u+\phi_{k})$, we evaluate $I_{k}=\int_{0}^{\infty}\frac{2w^{2}\sin(\omega_{k}u+\phi_{k})}{u(u^{2}+w^{2})}\,du$ by contour integration. (The lower limit in (7) is $\gamma_{1}^{(0)}\approx 17.8$, not 0; the difference $\int_{0}^{\gamma_{1}^{(0)}}$ is bounded and absorbed into the $O(c)$ term.)
Write $\frac{2w^{2}}{u(u^{2}+w^{2})}=\frac{2}{u}-\frac{2u}{u^{2}+w^{2}}$ and use the standard integrals (for $\omega>0$, $\operatorname{Re}(w)>0$):
$\int_{0}^{\infty}\frac{\sin(\omega u+\phi)}{u}\,du$ $=\frac{\pi}{2}\cos\phi+\text{(Ci terms)},$
$\int_{0}^{\infty}\frac{u\sin(\omega u+\phi)}{u^{2}+w^{2}}\,du$ $=\frac{\pi}{2}\operatorname{Re}\left[e^{i\phi}\,e^{-\omega w}\right]+\text{(Ci terms)}.$
The Ci (cosine integral) terms cancel in the difference. The key identity, verified by residue calculus at the poles $u=\pm iw$ of $1/(u^{2}+w^{2})$, is:
(8) $I_{k}=\pi\cos\phi_{k}\cdot\bigl(1-e^{-\omega_{k}\operatorname{Re}(w)}\cos(\omega_{k}\operatorname{Im}(w))\bigr)+\pi\sin\phi_{k}\cdot e^{-\omega_{k}\operatorname{Re}(w)}\sin(\omega_{k}\operatorname{Im}(w)).$
For $s=\sigma+it$: $\operatorname{Re}(w)=\sigma-1/2$, $\operatorname{Im}(w)=t$. The constant term ($1$ inside the parenthesis) gives a $t$-independent contribution: multiplied by $c/(\pi\alpha_{k}^{1/2})$ and summed, this produces $c\sum_{k}\cos(\phi_{k})/\alpha_{k}^{1/2}$, a convergent sum (by Lemma 2.1 with $T=0$) of magnitude $O(c)$. The oscillating terms combine to:
$I_{k}^{\mathrm{osc}}=\pi\,e^{-\omega_{k}(\sigma-1/2)}\bigl[-\cos\phi_{k}\cos(\omega_{k}t)+\sin\phi_{k}\sin(\omega_{k}t)\bigr]=-\pi\,\alpha_{k}^{-(\sigma-1/2)}\cos(\omega_{k}t+\phi_{k}),$
using $e^{-\omega_{k}(\sigma-1/2)}=\alpha_{k}^{-(\sigma-1/2)}$ and the angle-addition identity. Substituting into (7) with $E_{K}=-\frac{1}{\pi}\sum_{k=1}^{K}\alpha_{k}^{-1/2}\sin(\omega_{k}u+\phi_{k})$, the contribution of the $k$-th term is $c\,\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})$ (the factors of $\pi$ cancel). Summing over $k=1,\ldots,K$:
$\operatorname{Re}\log\frac{\Xi_{c,K}(s)}{\Xi_{0}(s)}=-c\sum_{k=1}^{K}\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c).$
The $O(c)$ error has two sources: the constant Fourier terms $c\sum_{k}\cos(\phi_{k})/\alpha_{k}^{1/2}$ (a convergent sum independent of $s$) and the lower-limit correction $c\int_{0}^{\gamma_{1}^{(0)}}(\cdots)\,du$ (bounded uniformly for $\sigma\geq 1$, $|t|\geq 2$). Both are independent of $K$; the total error is $c\cdot C_{0}+O(c^{2})$ where $C_{0}=C_{0}(\theta)$ is an explicit constant depending on the phase parameter $\theta$.
Taking $K\to\infty$: the LHS converges to $\operatorname{Re}\log(\Xi_{c}(s)/\Xi_{0}(s))$ (since $\Xi_{c,K}\to\Xi_{c}$ pointwise), and the trigonometric sum on the RHS converges for $\sigma\geq 1$ (absolutely for $\sigma>1$; conditionally at $\sigma=1$ by Lemma 2.1). The linearization error $\sum\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2}$ for the full series converges by Lemma 1.2, contributing $O(c^{2})$. The identity (6) follows with total error $O(c)$. ∎
Lemma 3.2.
$\log|Z_{0}(\sigma+it)|=O(1)$ uniformly for $\sigma\geq 1$ and $|t|\geq 2$.
Proof.
Since $N_{0}(T)=N_{\mathrm{main}}(T)$ (no error term), the Euler–Maclaurin formula converts the zero sum to an integral: $\log|\Xi_{0}(s)|=\sum_{n}\log|1+w^{2}/(\gamma_{n}^{(0)})^{2}|=\int_{\gamma_{1}^{(0)}}^{\infty}\log|1+w^{2}/u^{2}|\,N_{\mathrm{main}}^{\prime}(u)\,du+O(1)$. By Stirling’s formula, $\int\log|1+w^{2}/u^{2}|\,N_{\mathrm{main}}^{\prime}(u)\,du=\log|\Gamma_{\pi}(s)|+O(1)$ where $\Gamma_{\pi}(s)=\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)$ (the integral reproduces the Weierstrass product of $\Gamma$ at the Weyl density). Hence $\log|Z_{0}|=\log|\Xi_{0}/\Gamma_{\pi}|=O(1)$. Note that $Z_{0}$ has a simple pole at $s=1$ from $1/[\tfrac{1}{2}s(s-1)]$; the bound is for $|t|\geq 2$. ∎
Remark 3.3.
Consequently, $\log|Z_{c}(\sigma+it)|=-c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(1)$ for $\sigma\geq 1$, $|t|\geq 2$.
4. The Fourier-Analytic Mechanism
The bounded/unbounded transition has a simple Fourier-analytic explanation. The spectral identity (Proposition 3.1) gives $\log|Z_{c}/Z_{0}|(\sigma+it)=-c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c)$. For $\sigma>1$: the sum $\sum\alpha_{k}^{-\sigma}$ converges (since $\alpha_{k}=p_{k}$ and $\sum p^{-\sigma}<\infty$), so the triangle inequality gives $|\log|Z_{c}/Z_{0}||\leq c\sum\alpha_{k}^{-\sigma}+O(c)$—boundedness. At $\sigma=1$: the sum $\sum\alpha_{k}^{-1}$ diverges (Mertens [11, Thm. 2.7]), and the Dirichlet series structure of $Z_{c}/Z_{0}$ (Remark 6.1) enables a Phragmén–Lindelöf argument (§5) to transfer the divergence into unboundedness of $Z_{c}$ on $\sigma=1$. The transition is thus driven entirely by the marginal convergence/divergence of the damped Fourier amplitudes $\sum\alpha_{k}^{-\sigma}$ at $\sigma=1$.
5. Proof of the Main Theorem
Proof of Theorem 1.1.
The function $\Xi_{c}$ is the Hadamard product (2) with zeros $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$ defined via the full series $E$ as in (5). By Lemma 2.2, $\Xi_{c}$ is an entire function of order $1$.
Part (a): by construction, zeros lie at $s=1/2\pm i\gamma_{n}$. Part (b): $\Xi_{c}(1-s)=\Xi_{c}(s)$ (since $(1-s-1/2)^{2}=(s-1/2)^{2}$).
Part (c): the counting function $N_{c}(T)=\#\{n:\gamma_{n}\leq T\}$ satisfies $|N_{c}(T)-N_{\mathrm{main}}(T)|=O(\log T)$ because the shifts $\delta_{n}=O(c\gamma_{n}^{(0)}/\log\gamma_{n}^{(0)})$ (Lemma 2.2) can change $N_{c}(T)$ by at most $O(\log T)$ relative to $N_{\mathrm{main}}(T)$.
Part (d) (Boundedness for $\sigma>1$). By Lemma 1.2, the linearization error satisfies $\sum\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2}=O(c^{2})$. The spectral identity (Proposition 3.1) therefore holds for $\Xi_{c}$ with $O(c)$ error, giving
$\log|Z_{c}(\sigma+it)|\leq c\,\zeta_{\alpha}(\sigma)+O(1)<\infty$
for all $\sigma>1$, $|t|\geq 2$ (using $\log|Z_{0}|=O(1)$ from Remark 3.2).
Part (e) (Unboundedness on $\sigma=1$). The ratio $Z_{c}/Z_{0}$ is a Dirichlet series with generalized Euler product (Remark 6.1), absolutely convergent for $\sigma>1$, hence bounded on $\sigma=2$. We show it is unbounded on $\sigma=1$ by a Phragmén–Lindelöf argument (cf. [14, Thm. 8.4]).
Step 1. $Z_{c}/Z_{0}$ is unbounded in $\{\sigma>1,\;|t|\geq 2\}$. For any $N$, by Dirichlet’s approximation theorem there exists $t\leq(2\pi/\varepsilon)^{N}$ with $|\omega_{k}t+\phi_{k}-\pi|<\varepsilon\pmod{2\pi}$ for all $k\leq N$. At $\sigma=1+1/\log N$:
$\operatorname{Re}\log\frac{Z_{c}}{Z_{0}}(\sigma\!+\!it)\geq c\sum_{k=1}^{N}p_{k}^{-\sigma}-O(c)\geq\tfrac{c}{3}\log\log N-O(1),$
which grows without bound as $N\to\infty$ (since $\sum_{p\leq N}p^{-1-1/\log N}\geq e^{-2}\sum_{p\leq N}p^{-1}\sim e^{-2}\log\log N$ by Mertens).
Step 2. Suppose $|Z_{c}/Z_{0}(1+it)|\leq M$ for all $t$. By Phragmén–Lindelöf in the strip $1\leq\sigma\leq 2$ (the function has at most polynomial growth on $\sigma=1$ and is bounded on $\sigma=2$), $Z_{c}/Z_{0}$ would be bounded throughout the strip. This contradicts Step 1.
Therefore $Z_{c}/Z_{0}$ is unbounded on $\sigma=1$. Since $|Z_{0}(1+it)|=O(1)$ (Remark 3.2) and $|Z_{0}(1+it)|$ is bounded away from zero for $|t|\geq 2$, it follows that $\sup_{t}|Z_{c}(1+it)|=+\infty$.
The transition is thus: for $\sigma>1$, $|Z_{c}|\leq\exp(c\,\zeta_{\alpha}(\sigma)+O(1))$ (unconditional by Lemma 1.2); on $\sigma=1$, $\sup_{t}|Z_{c}(1+it)|=+\infty$ (unconditional by the Dirichlet–PLF argument above). (Note: $Z_{c}$ has a pole at $s=1$ from the factor $1/[\tfrac{1}{2}s(s-1)]$; the large values are for $|t|\to\infty$.) ∎
6. The Ratio $Z_{c}/Z_{0}$ and Dirichlet Series
Remark 6.1 (Dirichlet series structure).
Since $\alpha_{k}=p_{k}$, the spectral identity becomes
$\log\frac{Z_{c}}{Z_{0}}(s)=-c\sum_{k=1}^{\infty}e^{-i\phi_{k}}\,p_{k}^{-s}+O(c),$
which is an ordinary Dirichlet series supported on primes. Exponentiating:
(9) $\frac{Z_{c}}{Z_{0}}(s)=\prod_{p}\exp\!\bigl(-c\,e^{-i\phi_{k}}\,p^{-s}\bigr)=\sum_{n=1}^{\infty}a_{n}\,n^{-s},$
with $|a_{p}|=c$,   $|a_{p^{m}}|=c^{m}/m!$,   and $a_{mn}=a_{m}a_{n}$ for $\gcd(m,n)=1$. This converges absolutely for $\sigma>1$. Since both $\Xi_{c}$ and $\Xi_{0}$ satisfy $\Xi(s)=\Xi(1\!-\!s)$ with the same $\Gamma/\pi$ prefactor, the ratio inherits a functional equation:
$\frac{Z_{c}}{Z_{0}}(s)=\frac{Z_{c}}{Z_{0}}(1-s).$
The ratio $Z_{c}/Z_{0}$ is meromorphic: it has zeros at $\{1/2+i\gamma_{n}^{(c)}\}$ (from $Z_{c}$) and poles at $\{1/2+i\gamma_{n}^{(0)}\}$ (from $Z_{0}$), both on $\sigma=1/2$. It is a Dirichlet series (with generalized Euler product) that is bounded for $\sigma>1$ and unbounded on $\sigma=1$, with all its zeros provably on the critical line.
The construction separates the Euler product (which lives on the ratio) from the entireness (which lives on $\Xi_{c}$). Note that this variant uses $\alpha_{k}=p_{k}$ (rational frequencies) and produces a single function $\Xi_{c}$ with the same properties as in Theorem 1.1; the ratio $Z_{c}/Z_{0}$ is then a Dirichlet series, but $Z_{0}$ is not.
7. Almost-Periodicity
The transition at $\sigma=1$ has a natural interpretation in the hierarchy of almost-periodic function spaces.
Proposition 7.1.
Let $\sigma>\tfrac{1}{2}$.
  1. (i)
    For $\sigma>1$: the function $t\mapsto\log|Z_{c}/Z_{0}|(\sigma\!+\!it)$ is Bohr almost-periodic.
  2. (ii)
    For $\sigma>\tfrac{1}{2}$: the function $t\mapsto\log|Z_{c}/Z_{0}|(\sigma\!+\!it)$ is Besicovitch ($B^{2}$) almost-periodic with
    $\|{\log|Z_{c}/Z_{0}|}\|_{B^{2}}^{2}=\tfrac{1}{2}\,c^{2}\sum_{k=1}^{\infty}\alpha_{k}^{-2\sigma}<\infty.$
  3. (iii)
    The construction $E(T)$ is not $B^{2}$-AP: $\sum|a_{k}|^{2}=\frac{1}{\pi^{2}}\sum\alpha_{k}^{-1}=\infty$ (Mertens). The truncation $E_{K}\in B^{2}\text{-AP}$ with $\|E_{K}\|_{B^{2}}^{2}=\frac{1}{2\pi^{2}}\sum_{k=1}^{K}\alpha_{k}^{-1}$.
Proof.
(i) By Proposition 3.1, the function equals an absolutely convergent trigonometric series $c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c)$ for $\sigma>1$. Since $\sum\alpha_{k}^{-\sigma}$ converges for $\sigma>1$ [11, Thm. 2.7], the series defines a Bohr-AP function.
(ii) The Fourier–Bohr coefficients at exponents $\pm\omega_{k}$ have squared amplitudes $(c\,\alpha_{k}^{-\sigma}/2)^{2}$. Since $\sum\alpha_{k}^{-2\sigma}$ converges for $2\sigma>1$, the Riesz–Fischer theorem for $B^{2}$-AP functions gives the result.
(iii) The amplitudes of $E(T)$ are $a_{k}=1/(\pi\alpha_{k}^{1/2})$. Since $\sum|a_{k}|^{2}=(1/\pi^{2})\sum\alpha_{k}^{-1}\to\infty$ (by Mertens’ theorem [11, Thm. 2.7]), $E$ does not lie in $B^{2}\text{-AP}$ (its mean square diverges). For the truncation $E_{K}$, the finite sum confirms $E_{K}\in B^{2}\text{-AP}$. ∎
The AP hierarchy for $Z_{c}$ mirrors that of $\zeta$: Bohr-AP for $\sigma>1$ (Bohr [1]), $B^{2}$-AP for $\sigma>1/2$ (Bohr–Jessen [3]), with the passage from convergent to divergent Fourier norm at $\sigma=1$ being the “Bohr-AP to $B^{2}$-AP” transition. The transitions occur at the same critical $\sigma$ values because $\alpha_{k}\sim p_{k}$; the difference is the source: arithmetic (Euler product) for $\zeta$, purely analytic (spectral identity) for $Z_{c}$.
8. Observations
We illustrate the construction numerically with $c=0.05$ and a finite truncation $K=80$. (The computations use the shifted variant $\alpha_{k}=p_{k}+\theta$ with $\theta=e^{\sqrt{2}}$, which ensures global monotonicity of $F_{K}$ (Remark 2.3); the qualitative behavior is identical to the $\alpha_{k}=p_{k}$ case.)
8.1. First zeros of $Z_{c}$
The perturbed zeros $\gamma_{n}$ (defined by $F_{K}(\gamma_{n})=n$ for the truncation) and their shifts $\delta_{n}=\gamma_{n}-\gamma_{n}^{(0)}$:
$n$ $\gamma_{n}$ $\delta_{n}$ gap $E(\gamma_{n})$
1 17.9157 +0.068+0.068 0.011-0.011
2 23.2341 +0.062+0.062 5.318 0.013-0.013
3 27.7063 +0.035+0.035 4.472 0.008-0.008
4 31.6613 0.057-0.057 3.955 +0.015+0.015
5 35.4981 +0.030+0.030 3.837 0.008-0.008
10 51.7580 +0.024+0.024 3.061 0.008-0.008
20 78.6866 +0.014+0.014 2.522 0.006-0.006
50 144.5359 +0.005+0.005 2.038 0.003-0.003
100 236.8981 +0.045+0.045 1.808 0.026-0.026
Max shift: $|\delta_{n}|\leq 0.077$. Min gap: $1.67>0$. Gaps decrease as $\sim 2\pi/\log(\gamma_{n}/(2\pi))$, matching the Weyl law.
8.2. Comparison with $\zeta$ zero statistics
The microscopic statistics of $Z_{c}$’s zeros differ fundamentally from those of $\zeta$. With 1000 zeros computed ($t\in[17.9,1420.1]$):
Statistic $\zeta$ (conditional on RH) $Z_{c}$ Clock
RMS of error term $\sqrt{\frac{\log\log T}{2\pi^{2}}}\to\infty$ $0.013$ (constant) 0
Spacing std (normalized) $\approx 0.42$ (GUE) $0.020$ 0
Pair correlation $1-\mathrm{sinc}^{2}\alpha$ peaks at integers $\delta$’s
Number variance $\Sigma^{2}(L)$ $\sim\tfrac{2}{\pi^{2}}\log L$ $\lesssim 0.25$ 0
Distribution of $S$ / $E$ Gaussian bounded, non-Gaussian 0
The zeros of $Z_{c}$ are super-rigid (nearly crystalline), with normalized spacing std two orders of magnitude below GUE. The construction captures the macroscopic transition (bounded/unbounded, AP hierarchy) but not the microscopic GUE statistics, which arise from the arithmetic content (Euler product, prime correlations) that $Z_{c}$ deliberately lacks.
8.3. Value distribution
Figure 1 shows the complex-plane images of $\log(Z_{c}/Z_{0})(\sigma+it)$, $\log\zeta(\sigma+it)$, and $\log L(\sigma+it,\chi_{4})$ as $t$ ranges over $[20,300]$, for $\sigma=1.25,1.0,0.75$.
The analytic formula gives
(10) $\log\frac{Z_{c}}{Z_{0}}(s)=-c\sum_{k=1}^{\infty}\alpha_{k}^{-s}\,e^{i\phi_{k}}+O(c),$
which has equal real and imaginary RMS: $\langle(\operatorname{Re}\log\frac{Z_{c}}{Z_{0}})^{2}\rangle=\langle(\operatorname{Im}\log\frac{Z_{c}}{Z_{0}})^{2}\rangle=\tfrac{c^{2}}{2}\sum\alpha_{k}^{-2\sigma}$. The value distribution is isotropic in the complex plane, as for $\log\zeta$ via Bohr–Jessen [3].
The quantitative difference is dramatic: at $\sigma=0.75$, the RMS of $\log(Z_{c}/Z_{0})$ is approximately $0.046$, while $\log\zeta$ has RMS $\approx 0.88$. The construction’s value distribution is confined to a bounded disk, whereas $|\log\zeta(\sigma+it)|$ is unbounded. This means $Z_{c}/Z_{0}$ cannot satisfy Voronin universality [15], and its moments do not exhibit the $(\log T)^{k^{2}}$ growth of $\zeta$. The value distribution is isotropic but concentrated—another manifestation of the missing arithmetic.
Figure 1. Complex-plane image of $\log f(\sigma+it)$ for $t\in[20,300]$. Left: $\log(Z_{c}/Z_{0})$; center: $\log\zeta$; right: $\log L(s,\chi_{4})$. Rows: $\sigma=1.25,1.0,0.75$. Color encodes $t$ (yellow==early, purple==late). Note the ${\sim}20{\times}$ difference in scale between the left column and the others.
Figure 2. Shape comparison: $\log(Z_{c}/Z_{0})$ rescaled to match the RMS of $\log\zeta$ (left) vs. $\log\zeta$ (right). The rescaled $Z_{c}$ fills a banded region (reflecting the coherent trigonometric structure) rather than the isotropic spirals of $\log\zeta$ (reflecting multiplicative independence of primes).
8.4. Numerical evidence for the linearization condition
Lemma 1.2 establishes that $\sum_{n=1}^{\infty}E(\gamma_{n}^{(0)})^{2}/n^{2}<\infty$. We verify this numerically with $K=80$ terms and $N=5000$ zero positions.
Growth of $|E(\gamma_{n}^{(0)})|$. The values $|E(\gamma_{n}^{(0)})|$ remain bounded: over the first 5000 zeros, $\max_{n}|E(\gamma_{n}^{(0)})|\approx 0.813$. A power-law regression gives $|E(\gamma_{n}^{(0)})|\sim n^{-0.025}$ (essentially flat), and a logarithmic fit yields $|E(\gamma_{n}^{(0)})|\sim(\log n)^{-0.16}$ (very slightly decreasing). The Abel-summation bound $|E(T)|=O(T)$ grossly overestimates the actual behavior at zero positions.
Mean square and Weyl equidistribution. The empirical mean square agrees with the Weyl prediction:
NN $\frac{1}{N}\sum_{n=1}^{N}E(\gamma_{n}^{(0)})^{2}$ $\frac{1}{2\pi^{2}}\sum_{k=1}^{K}\alpha_{k}^{-1}$ ratio
100 0.0688 0.0654 1.053
500 0.0668 0.0654 1.022
1000 0.0649 0.0654 0.994
5000 0.0648 0.0654 0.991
The ratio stabilizes near $1$, consistent with joint equidistribution of the $K$ phases $\{\omega_{k}\gamma_{n}^{(0)}+\phi_{k}\}$ modulo $2\pi$ (cross terms vanish, diagonal terms give the Mertens sum).
Convergence of $\sum E^{2}/n^{2}$. The partial sums of $\sum_{n=1}^{N}E(\gamma_{n}^{(0)})^{2}/n^{2}$ stabilize rapidly:
NN $\sum_{n=1}^{N}E^{2}/n^{2}$ $\sum_{n=1}^{N}\delta_{n}^{2}/\gamma_{n}^{2}$ last term $E^{2}/n^{2}$
100 0.0891 $3.79\times 10^{-5}$ $2.80\times 10^{-5}$
1000 0.0897 $88\times 10^{-5} 3.3$$30\times 10^{-7}$
5000 0.0898 $3.89\times 10^{-5}$ $3.00\times 10^{-11}$
The growth from $N=1000$ to $N=5000$ is only $0.06\%$. The linearization error $\sum\delta_{n}^{2}/\gamma_{n}^{2}\approx 3.9\times 10^{-5}$ is far below $c^{2}=2.5\times 10^{-3}$.
Spectral identity remainder. The remainder $R(t)$ in the spectral identity $\log|\Xi_{c}/\Xi_{0}|(\sigma\!+\!it)=-c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+R(t)$ has:
$\sigma$ $\operatorname{mean}(R)$ $\operatorname{std}(R)$ $\max|R|$
1.50 0.015-0.015 $0.021$ $0.081$
1.10 0.016-0.016 $0.030$ $0.122$
1.01 0.016-0.016 $0.034$ $0.136$
1.00 0.016-0.016 $0.034$ $0.138$
The remainder is $O(c)$ as claimed in Proposition 3.1, and the contribution from linearization ($\sim 10^{-5}$) is negligible compared to the constant and lower-limit terms.
9. Open Problems
Several directions for further investigation suggest themselves.
GUE statistics. The zeros of $Z_{c}$ are super-rigid (nearly crystalline spacing, §8.2), far from the GUE statistics of $\zeta$. What additional structure on the zero perturbation $E$ would force GUE pair correlation onto the zeros of $\Xi_{c}$?
Pointwise bounds on $E$. Numerically, $|E(\gamma_{n}^{(0)})|\leq 1.05$ for all computed nn8.4), vastly smaller than the Abel bound $O(\gamma_{n})$. Proving $|E(\gamma_{n}^{(0)})|=O(1)$ would give a much tighter linearization estimate than Lemma 1.2.
Growth rate on $\sigma=1$. The Phragmén–Lindelöf argument gives $\sup_{t}|Z_{c}(1+it)|=+\infty$ but not the rate. For $\zeta$, $|\zeta(1+it)|=\Omega(\log\log t)$ [14, Thm. 8.5]. Does $|Z_{c}(1+it)|$ achieve the same rate?
10. Concluding Remarks
The function $\Xi_{c}$ provides the first explicit entire function of order $1$ that provably satisfies an analogue of the Riemann Hypothesis (all zeros on $\sigma=1/2$) together with the sharp bounded/unbounded transition at $\sigma=1$ characteristic of $\zeta$. The construction demonstrates that these properties are simultaneously achievable and provable, given sufficient control over the zero distribution.
The ratio $Z_{c}/Z_{0}$ is a Dirichlet series with generalized Euler product, providing a concrete object whose zeros are provably on the critical line and whose analytic properties (boundedness, value distribution, almost-periodicity) can be studied unconditionally. In the context of the Selberg class [12, 9], the construction shows that an RH-like zero distribution is compatible with the $\sigma=1$ transition—the two properties reinforce rather than obstruct each other.
The microscopic statistics of $Z_{c}$’s zeros differ from those of $\zeta$8.2): the zeros are nearly crystalline (super-rigid spacing) rather than GUE-distributed, and $Z_{c}$ does not satisfy Voronin universality. These differences reflect the specific arithmetic content—prime correlations, multiplicative independence of $p^{-it}$—that $\zeta$ possesses and $Z_{c}$ deliberately lacks. Understanding which additional structure would force GUE statistics onto the zeros of an RH-satisfying function remains an interesting open problem.