### v2

# An Explicit Entire Function of Order One with All Zeros on a Line and Bounded in a Half-Plane

Abstract.
We construct a single explicit entire function $\Xi_{c}(s)$ of order 1, with all zeros provably on $\operatorname{Re}(s)=1/2$, satisfying a functional equation $\Xi_{c}(s)=\Xi_{c}(1-s)$, whose normalized form $Z_{c}(s)=\Xi_{c}(s)/[\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)]$ is uniformly bounded for $\operatorname{Re}(s)>1+\delta$ (any $\delta>0$) yet satisfies $\sup_{t}|Z_{c}(1+it)|=+\infty$. The function thus satisfies an analogue of the Riemann Hypothesis together with the sharp bounded/unbounded transition at $\sigma=1$ characteristic of $\zeta$. The transition is controlled by a Dirichlet series $D(s)=\sum e^{-ik\theta}\,p_{k}^{-s}$ whose absolute convergence for $\sigma>1$ and divergence at $\sigma=1$ drive the dichotomy, in direct analogy with $\log\zeta(s)=\sum p^{-s}+O(1)$. The key technical input is a dyadic large-sieve estimate establishing the linearization condition that connects the Hadamard product to $D$. The construction and proofs were developed in collaboration with Claude (Anthropic); see Acknowledgments.
1. Introduction
The Riemann zeta function exhibits a sharp transition at $\sigma=1$:
  • For $\sigma>1$: $|\zeta(\sigma+it)|\leq\zeta(\sigma)<\infty$.
  • On $\sigma=1$: $\limsup_{t\to\infty}|\zeta(1+it)|=\infty$, with $|\zeta(1+it)|=\Omega(\log\log t)$ [17, Thm. 8.5].
The Selberg class [15, 10] axiomatizes $L$-functions via five properties: Dirichlet series, analytic continuation, functional equation, Ramanujan bound, and Euler product. These axioms are top-down: they prescribe what $L$-functions should look like, based on the structure of $\zeta$ and its generalizations. But they do not reveal which properties are truly necessary for the Riemann Hypothesis or the $\sigma=1$ transition. The Euler product is widely regarded as essential [7, 16]: without it, the Davenport–Heilbronn function [4] has zeros off the critical line, and Bombieri–Garrett’s pseudo-Laplacians [2] fail to capture all zeros spectrally. The de Branges space framework [5] provides a natural setting for entire functions with real zeros, but verifying the required positivity conditions is equivalent to RH itself. For general background on $L$-functions and the analytic tools used here, see [9].
A natural question is whether one can construct a function that provably satisfies an analogue of RH together with the $\sigma=1$ transition. For $\zeta$ itself, the location of zeros is the central open problem; but for a suitably designed function, one can hope to build in the zeros-on-the-line property while retaining enough structure for the $\sigma=1$ transition to hold.
The mechanism behind $\zeta$’s transition is well understood at the level of Dirichlet series: $\log\zeta(s)=P(s)+O(1)$ for $\sigma>1$, where $P(s)=\sum p^{-s}$ is the prime zeta function [17, §1.6, §3.2]. Boundedness for $\sigma>1$ follows from the absolute convergence of $\sum p^{-\sigma}$; unboundedness on $\sigma=1$ from its divergence (Mertens [14, Thm. 2.7]). The zero-counting error $S(T)=\pi^{-1}\arg\zeta(\tfrac{1}{2}+iT)$ satisfies $S(T)\approx-\pi^{-1}\sum_{p}\sin(T\log p)/\sqrt{p}$ in $L^{2}$ [17, Thm. 14.22]—the imaginary part of $P(\tfrac{1}{2}+iT)/\pi$.
Our construction replaces $P(s)$ with a twisted variant $D(s)=\sum e^{-ik\theta}\,p_{k}^{-s}$ (with $\theta\in(0,2\pi)$), uses its restriction to $\sigma=1/2$ to perturb the zeros of a Hadamard product, and proves that the resulting function’s log-ratio with the unperturbed product is approximately cD(s)-cD(s)—the “spectral identity” of this paper.
Theorem 1.1 (Main Theorem).
Fix $\theta\in(0,2\pi)$ (e.g., $\theta=1$) and $c>0$ sufficiently small. Define $\Xi_{c}(s)$ as in §2, with $Z_{c}(s)=\Xi_{c}(s)/[\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)]$. Then:
  1. (a)
    All zeros of $\Xi_{c}$ lie on $\operatorname{Re}(s)=1/2$.
  2. (b)
    $\Xi_{c}(s)=\Xi_{c}(1-s)$.
  3. (c)
    The zero-counting function satisfies $N_{c}(T)=\frac{T}{2\pi}\log\frac{T}{2\pi}-\frac{T}{2\pi}+O(\log T)$.
  4. (d)
    For each $\delta>0$, $\sup_{t}|Z_{c}(1+\delta+it)|\leq\exp\bigl(c\,P(1+\delta)+O(1)\bigr)$, where $P(\sigma)=\sum_{p}p^{-\sigma}$. Moreover $P(\sigma)\to 0$ as $\sigma\to\infty$, so the bound tends to $e^{O(1)}$: the ratio $\Xi_{c}/\Xi_{0}$ is uniformly close to a constant for $\sigma$ large.
  5. (e)
    $\sup_{t}|Z_{c}(1+it)|=+\infty$.
  6. (f)
    The ratio $\Xi_{c}/\Xi_{0}$ admits the approximate Euler product $\Xi_{c}(s)/\Xi_{0}(s)=\kappa_{0}\cdot\prod_{k}\exp(-c\,e^{-ik\theta}p_{k}^{-s})\cdot M(s)$ for $\sigma>1$, where $\kappa_{0}\neq 0$ is an explicit constant, $M(s)$ is bounded (with $M\to 1$ as $c\to 0$), and the Euler product is itself a Dirichlet series with multiplicative coefficients.
2. The Dirichlet Series and Construction
2.1. The Dirichlet series $D(s)$
Fix $\theta\in(0,2\pi)$ (e.g., $\theta=1$), an amplitude $c>0$, and let $p_{1}<p_{2}<\cdots$ denote the primes. Define the Dirichlet series
(1) $D(s)=\sum_{k=1}^{\infty}e^{-ik\theta}\,p_{k}^{-s},$
which converges absolutely for $\sigma>1$ (since $\sum p_{k}^{-\sigma}<\infty$) and conditionally for $\sigma>0$, by Abel summation: the partial sums $G_{K}=\sum_{k=1}^{K}e^{-ik\theta}$ satisfy $|G_{K}|\leq 2/|1-e^{-i\theta}|=:C_{\theta}$ (finite since $\theta\in(0,2\pi)$), and the amplitudes $p_{k}^{-\sigma}$ decrease monotonically to zero. Viewing $D$ as an ordinary Dirichlet series $\sum b_{n}\,n^{-s}$ (with $b_{n}=e^{-ik\theta}$ for $n=p_{k}$, zero otherwise), the partial sums $B(x)=\sum_{n\leq x}b_{n}$ satisfy $|B(x)|\leq C_{\theta}$, so the abscissa of convergence is $\sigma_{c}=0$ (cf. [14, Thm. 1.3]). In the strip of conditional convergence, [14, Thm. 1.5] gives
(2) $|D(\sigma+it)|\ll(|t|+4)^{1-\sigma+\varepsilon}$
for any fixed $\sigma\in(0,1)$ and $\varepsilon>0$. For $\sigma>1$ the stronger bound $|D(\sigma+it)|\leq P(\sigma)=\sum p^{-\sigma}$ holds by absolute convergence.
Define the perturbation function
(3) $E(T)=\frac{1}{\pi}\,\operatorname{Im}\bigl[D(\tfrac{1}{2}+iT)\bigr]=-\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(T\log p_{k}+\phi_{k})}{p_{k}^{1/2}},$
where $\phi_{k}=k\theta$. The convergence of $E(T)$ for each $T>0$ follows from that of $D(1/2+iT)$, and by (2), $|E(T)|=O(|T|^{1/2+\varepsilon})$. We write $\omega_{k}=\log p_{k}$ for brevity in Fourier-analytic arguments.
2.2. The perturbed zeros
Define the smooth counting function
(4) $N_{\mathrm{main}}(T)=\frac{T}{2\pi}\log\frac{T}{2\pi}-\frac{T}{2\pi}+\frac{7}{8},\qquad N_{\mathrm{main}}^{\prime}(T)=\frac{1}{2\pi}\log\frac{T}{2\pi}.$
Let $\{\gamma_{n}^{(0)}\}$ be the “regular” sequence defined by $N_{\mathrm{main}}(\gamma_{n}^{(0)})=n$, with Hadamard product $\Xi_{0}(s)$. Define the “perturbed” sequence $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$ with
(5) $\delta_{n}=-\frac{cE(\gamma_{n}^{(0)})}{N_{\mathrm{main}}^{\prime}(\gamma_{n}^{(0)})},$
and let $\Xi_{c}(s)$ be the Hadamard product (cf. [6, §1.10])
(6) $\Xi_{c}(s)=\prod_{n=1}^{\infty}\left(1+\frac{(s-\tfrac{1}{2})^{2}}{\gamma_{n}^{2}}\right).$
Lemma 2.1 (Well-definedness).
For $c$ sufficiently small, $\gamma_{n}>0$ for all $n$, $\sum\gamma_{n}^{-2}<\infty$, and $\Xi_{c}(s)$ is an entire function of order $1$.
Proof.
Since $|E(T)|=O(|T|^{1/2+\varepsilon})$, $|\delta_{n}|=O(c(\gamma_{n}^{(0)})^{1/2+\varepsilon}/\log\gamma_{n}^{(0)})=o(\gamma_{n}^{(0)})$, so $\gamma_{n}>\gamma_{n}^{(0)}/2$ for all $n$ sufficiently large. For the finitely many small $n$, $\delta_{n}$ is finite (since $E(\gamma_{n}^{(0)})$ converges for each $n$) and proportional to $c$, so taking $c$ small enough ensures $\gamma_{n}>0$ for all $n$. Since $\gamma_{n}^{(0)}\sim 2\pi n/\log n$, $\sum\gamma_{n}^{-2}<\infty$ and the exponent of convergence equals $1$. ∎
We set $Z_{c}(s)=\Xi_{c}(s)/\Gamma_{\pi}(s)$ and $Z_{0}(s)=\Xi_{0}(s)/\Gamma_{\pi}(s)$, where $\Gamma_{\pi}(s)=\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)$. Note that $Z_{c}/Z_{0}=\Xi_{c}/\Xi_{0}$ (the $\Gamma_{\pi}$ factors cancel).
3. The Spectral Identity
The central result relates the Hadamard product ratio $\Xi_{c}/\Xi_{0}$ to the Dirichlet series $D$.
Proposition 3.1 (Spectral Identity).
Let $\omega_{k}=\log p_{k}$ and $\phi_{k}=k\theta$ as in §2. For $\sigma>1/2$ and $|t|\geq 2$:
(7) $\operatorname{Re}\log\frac{\Xi_{c}}{\Xi_{0}}(\sigma+it)=-c\sum_{k}p_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(1).$
Equivalently, there exists a bounded analytic function $R(s)$ with $R(s)=O(1)$ uniformly for $\sigma>1/2$, $|t|\geq 2$, such that
(8) $\log\frac{\Xi_{c}}{\Xi_{0}}(s)=-c\,D(s)+R(s).$
The cosine sum converges absolutely for $\sigma>1$ and conditionally for $\sigma>0$ (as $\operatorname{Re}\,D(\sigma+it)$).
Proof.
We establish (7) first, then deduce (8) by analyticity. Throughout, $c$ is fixed, so all implicit constants may depend on $c$ and $\theta$.
Step 1 (Linearization). By (5), $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$. Each Hadamard factor satisfies:
$\log\frac{1+(s-\tfrac{1}{2})^{2}/\gamma_{n}^{2}}{1+(s-\tfrac{1}{2})^{2}/(\gamma_{n}^{(0)})^{2}}=\frac{-2w^{2}\delta_{n}}{\gamma_{n}^{(0)}((\gamma_{n}^{(0)})^{2}+w^{2})}+O(\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2})$
where w=s1/2w=s-1/2.
Step 2 (Sum to integral). For each $K$, let $\Xi_{c,K}$ be the Hadamard product with zeros perturbed using $E_{K}$ (first $K$ terms of $E$). Since $|E_{K}|\leq\frac{1}{\pi}\sum_{k=1}^{K}p_{k}^{-1/2}$ (a constant for each $K$), the perturbations are $\delta_{n,K}=O(1/\log\gamma_{n}^{(0)})$, giving convergent quadratic remainder $\sum\delta_{n,K}^{2}/(\gamma_{n}^{(0)})^{2}<\infty$. The linearized sum is a Riemann sum for $\int\frac{2w^{2}\,cE_{K}(u)}{u(u^{2}+w^{2})}\,du$ (one term per zero, spacing $\sim 1/N_{\mathrm{main}}^{\prime}$). The integrand is smooth for $u>0$ and $\sigma>1/2$ (the poles of $1/(u^{2}+w^{2})$ lie off the real axis), so by Euler–Maclaurin:
(9) $\operatorname{Re}\log\frac{\Xi_{c,K}(s)}{\Xi_{0}(s)}=-c\int_{\gamma_{1}^{(0)}}^{\infty}\operatorname{Re}\frac{2w^{2}\,E_{K}(u)}{u(u^{2}+w^{2})}\,du+O(1).$
Step 3 (Evaluation of the integral). Using the partial fraction $\frac{2w^{2}}{u(u^{2}+w^{2})}=\frac{2}{u}-\frac{2u}{u^{2}+w^{2}}$ and the standard integrals $\int_{0}^{\infty}\frac{\sin\omega u}{u}\,du=\frac{\pi}{2}$ (Dirichlet) and $\int_{0}^{\infty}\frac{u\sin\omega u}{u^{2}+\beta^{2}}\,du=\frac{\pi}{2}e^{-\omega\beta}$ for $\omega>0$, $\operatorname{Re}\beta>0$ ([8, §3.723], extended to complex $\beta$ by analytic continuation), the contribution of each component $\sin(\omega_{k}u+\phi_{k})$ decomposes into convergent pieces. [^1] The individual integrals $\int\sin(\omega u+\phi)/u\,du$ and $\int u\sin(\omega u+\phi)/(u^{2}+w^{2})\,du$ each diverge at $u=0$ when $\sin\phi\neq 0$ (from the $\cos(\omega u)/u$ term), but the divergent parts cancel in the partial-fraction difference. The oscillatory contribution of the $k$-th component is $\pi\,p_{k}^{-(\sigma-1/2)}\cos(\omega_{k}t+\phi_{k})$ (using $e^{-\omega_{k}(\sigma-1/2)}=p_{k}^{-(\sigma-1/2)}$ and the angle-addition identity); the non-oscillatory part $\pi\cos\phi_{k}(1-p_{k}^{-(\sigma-1/2)})$ sums to a bounded quantity (convergent by Abel summation, depending on $\sigma$ but not on $t$).
Substituting into (9) with the amplitudes $1/(\pi p_{k}^{1/2})$ from $E_{K}$, the $k$-th term contributes $c\,p_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})$. Including the bounded lower-limit correction $\int_{0}^{\gamma_{1}^{(0)}}(\cdots)\,du$:
(10) $\operatorname{Re}\log\frac{\Xi_{c,K}(s)}{\Xi_{0}(s)}=-c\sum_{k=1}^{K}p_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(1)+O\Bigl(\sum_{n}\delta_{n,K}^{2}/(\gamma_{n}^{(0)})^{2}\Bigr),$
where the $O(1)$ accounts for the non-oscillatory Fourier terms, the lower-limit correction, and the Euler–Maclaurin error (all independent of $K$), while the final term is the quadratic Taylor remainder from Step 1. For each fixed $K$ this remainder is finite, but it grows with $K$ (since $M_{K}$ grows).
Taking $K\to\infty$: the cosine sum converges (absolutely for $\sigma>1$; conditionally for $\sigma>0$), and $\Xi_{c,K}\to\Xi_{c}$ (since the zero sequences converge). The only obstruction to passing (10) to the limit is the Taylor remainder: for the full (non-truncated) series, $\delta_{n}=-cE(\gamma_{n}^{(0)})/N_{\mathrm{main}}^{\prime}(\gamma_{n}^{(0)})$ and $|E(\gamma_{n}^{(0)})|$ is no longer bounded by a fixed constant (its RMS grows like $\sqrt{\log\log n}$ by (14)). The quadratic remainder $\sum\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2}$ converges iff $\sum E(\gamma_{n}^{(0)})^{2}/n^{2}<\infty$, which is the content of Lemma 3.2 below. This establishes (7).
Step 4 (Complex extension). Equation (7) controls the real part of $\log(\Xi_{c}/\Xi_{0})+cD$; we now upgrade to the full complex identity. Define $F(s)=\log(\Xi_{c}(s)/\Xi_{0}(s))+cD(s)$. Both terms are analytic for $\sigma>1/2$: the Hadamard ratio $\Xi_{c}/\Xi_{0}$ is nonvanishing there (all zeros of $\Xi_{0}$ and $\Xi_{c}$ lie on $\sigma=1/2$), and $D$ converges for $\sigma>0$. By (7) (proved above without reference to (8)), $\operatorname{Re}F(s)=O(1)$ for $\sigma>1/2$, $|t|\geq 2$. Evaluating at $s=2$: $D(2)$ converges absolutely, and $\log(\Xi_{c}(2)/\Xi_{0}(2))$ is finite (each Hadamard factor at $s=2$ differs from $1$ by $O(\delta_{n}/\gamma_{n}^{2})$, and $\sum|\delta_{n}|/\gamma_{n}^{2}<\infty$), so $F(2)=O(1)$. Since $F$ is analytic with bounded real part, the Borel–Carathéodory theorem gives $|F(s)|=O(1)$ for $\sigma>1/2$. Setting $R(s)=F(s)$ gives (8). ∎
3.1. Linearization convergence
The following lemma ensures that the Taylor linearization in Step 1 has convergent remainder for the full (non-truncated) series, which is the key input for passing the spectral identity from finite truncations to $\Xi_{c}$ itself.
Lemma 3.2.
(11) $\sum_{n=1}^{\infty}\frac{E(\gamma_{n}^{(0)})^{2}}{n^{2}}<\infty.$
Proof.
Decompose $\mathbb{N}$ into dyadic blocks $I_{j}=[2^{j},2^{j+1})$ for $j\geq 1$. It suffices to show $\sum_{j}S_{j}<\infty$ where $S_{j}=\sum_{n\in I_{j}}E(\gamma_{n}^{(0)})^{2}/n^{2}$. For each $j$, set $K_{j}=\lceil 2^{1.1\,j}\rceil$ and split $E=E_{K_{j}}+R_{K_{j}}$, where $E_{K_{j}}$ is the first $K_{j}$ terms. By $(a+b)^{2}\leq 2a^{2}+2b^{2}$: $S_{j}\leq 2(\text{HEAD}_{j}+\text{TAIL}_{j})$.
Head (large sieve). $E_{K_{j}}$ is a trigonometric polynomial with $K_{j}$ terms and frequency range $\Lambda_{j}=\omega_{K_{j}}-\omega_{1}=O(j)$. The zeros $\gamma_{n}^{(0)}$ have minimum spacing $\Delta\geq c_{1}/j$ for $n\leq 2^{j+1}$. By the large sieve inequality for trigonometric polynomials at well-spaced points (cf. [13, Thm. 1], [14, §7]):
$\sum_{n\leq 2^{j+1}}\!|E_{K_{j}}(\gamma_{n}^{(0)})|^{2}\leq(\pi/\Delta+\Lambda_{j})\,\tfrac{1}{\pi^{2}}\sum_{k\leq K_{j}}p_{k}^{-1}\leq C_{1}\,j\log j,$
since $\sum_{k\leq K_{j}}p_{k}^{-1}\sim\log\log p_{K_{j}}\sim\log j$ (Mertens). Therefore $\text{HEAD}_{j}\leq C_{1}\,j\log j\,/\,4^{j}$, and $\sum_{j}j\log j/4^{j}<\infty$.
Tail (Abel summation). The tail $\sum_{k>K}(\omega_{k+1}-\omega_{k})/p_{k}^{1/2}$ is a left Riemann sum for the decreasing function $e^{-\omega/2}$. By Bertrand’s postulate (mesh width $\leq\log 2$): $\sum_{k>K}(\omega_{k+1}-\omega_{k})/p_{k}^{1/2}\leq\sqrt{2}\int_{\omega_{K}}^{\infty}e^{-\omega/2}\,d\omega=2\sqrt{2}\,p_{K}^{-1/2}$. Hence $|R_{K_{j}}(\gamma_{n}^{(0)})|\leq C_{2}\,2^{j(1-\varepsilon)/2}/j^{3/2}$ (with $\varepsilon=0.1$, using $\gamma_{n}\leq C_{3}\,2^{j}/j$ and $\sqrt{p_{K_{j}}}\geq C_{4}\,2^{(1+\varepsilon)j/2}\sqrt{j}$). Therefore $\text{TAIL}_{j}\leq C_{5}/(j^{3}\cdot 2^{\varepsilon j})$, and $\sum_{j}1/(j^{3}\cdot 2^{\varepsilon j})<\infty$.
Both contributions decay exponentially, so $\sum S_{j}<\infty$. ∎
3.2. The base function $Z_{0}$
Lemma 3.3.
$\log|Z_{0}(\sigma+it)|=O(1)$ uniformly for $\sigma\geq 1$ and $|t|\geq 2$. In particular, $|Z_{0}|$ is both bounded and bounded away from zero in this region.
Proof.
By construction, $\Xi_{0}$ has zero-counting function $N_{0}(T)=N_{\mathrm{main}}(T)$ (no error term). By Euler–Maclaurin, $\log|\Xi_{0}(s)|=\int_{\gamma_{1}^{(0)}}^{\infty}\log|1+w^{2}/u^{2}|\,N_{\mathrm{main}}^{\prime}(u)\,du+O(1)$. Stirling’s formula identifies this integral with $\log|\Gamma_{\pi}(s)|+O(1)$, giving $\log|Z_{0}|=\log|\Xi_{0}/\Gamma_{\pi}|=O(1)$. The second claim follows since $|\!\log|Z_{0}|\!|\leq C$ implies $e^{-C}\leq|Z_{0}|\leq e^{C}$. ∎
4. Proof of the Main Theorem
Proof of Theorem 1.1.
Parts (a)–(c) hold by construction; (d) follows from the spectral identity (Proposition 3.1) with linearization error controlled by Lemma 3.2; and (e) via a Phragmén–Lindelöf argument below.
The function $\Xi_{c}$ is the Hadamard product (6) with zeros $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$ defined via $E$ as in (5). By Lemma 2.1, $\Xi_{c}$ is an entire function of order $1$.
Part (a): by construction, zeros lie at $s=1/2\pm i\gamma_{n}$. Part (b): $\Xi_{c}(1-s)=\Xi_{c}(s)$ (since $(1-s-1/2)^{2}=(s-1/2)^{2}$).
Part (c): the counting function $N_{c}(T)=\#\{n:\gamma_{n}\leq T\}$ satisfies $|N_{c}(T)-N_{\mathrm{main}}(T)|=O(\log T)$. Indeed, zeros near height $T$ have index $n\sim T\log T/(2\pi)$ and perturbation $|\delta_{n}|=O(T^{1/2+\varepsilon}/\log T)$ by (2). In an interval of this length around $T$, $N_{\mathrm{main}}$ changes by $O(N_{\mathrm{main}}^{\prime}(T)\cdot T^{1/2+\varepsilon}/\log T)=O(T^{1/2+\varepsilon}/\log^{2}T)=o(\log T)$.
Part (d) (Boundedness for $\sigma>1$). The spectral identity (Proposition 3.1) gives
$\log|Z_{c}(\sigma+it)|=\operatorname{Re}\log\frac{\Xi_{c}}{\Xi_{0}}(s)+\log|Z_{0}(s)|\leq c\,P(\sigma)+O(1)$
for all $\sigma>1$, $|t|\geq 2$ (using $|\sum p_{k}^{-\sigma}\cos(\cdots)|\leq P(\sigma)$ and Lemma 3.3).
Part (e) (Unboundedness on $\sigma=1$). We show $\sup_{t}|\Xi_{c}/\Xi_{0}(1+it)|=+\infty$ by a Phragmén–Lindelöf argument (cf. [17, Thm. 8.4]).
Step 1. $\Xi_{c}/\Xi_{0}$ is unbounded in $\{\sigma>1,\;|t|\geq 2\}$. For any $N$, by Dirichlet’s approximation theorem there exists $t\leq(2\pi/\varepsilon)^{N}$ with $|\omega_{k}t+\phi_{k}-\pi|<\varepsilon\pmod{2\pi}$ for all $k\leq N$. At $\sigma=1+1/\log N$:
$\operatorname{Re}\log\frac{\Xi_{c}}{\Xi_{0}}(\sigma\!+\!it)\geq c\sum_{k=1}^{N}p_{k}^{-\sigma}-O(c)\geq\tfrac{c}{e}\log\log N-O(1),$
which grows without bound as $N\to\infty$ (since $p_{k}^{-1/\log N}=e^{-\log p_{k}/\log N}\geq e^{-1}$ for $p_{k}\leq N$, so $\sum_{p\leq N}p^{-1-1/\log N}\geq e^{-1}\sum_{p\leq N}p^{-1}\sim e^{-1}\log\log N$).
Step 2. Suppose $|\Xi_{c}/\Xi_{0}(1+it)|\leq M$ for all $t$. Since $\Xi_{c}$ and $\Xi_{0}$ are entire of order $1$, $|\Xi_{c}(\sigma+it)/\Xi_{0}(\sigma+it)|\leq\exp(C|t|^{1+\varepsilon})$ for any $\varepsilon>0$ (the ratio is analytic in $\sigma\geq 1$ since all zeros of $\Xi_{0}$ lie on $\sigma=1/2$). The Phragmén–Lindelöf principle for the strip $1\leq\sigma\leq 2$—bounded on $\sigma=1$ (by assumption) and on $\sigma=2$ (by part (d)), with at most exponential growth—gives boundedness throughout. This contradicts Step 1.
Therefore $\Xi_{c}/\Xi_{0}$ is unbounded on $\sigma=1$, and since $Z_{c}/Z_{0}=\Xi_{c}/\Xi_{0}$ and $|Z_{0}(1+it)|$ is bounded away from zero for $|t|\geq 2$ (Lemma 3.3), $\sup_{t}|Z_{c}(1+it)|=+\infty$.
Part (f) (Approximate Euler product). This is Proposition 5.1 below. ∎
5. The Ratio $\Xi_{c}/\Xi_{0}$ and Dirichlet Series
Proposition 5.1.
For $\sigma>1$, the ratio $\Xi_{c}/\Xi_{0}$ admits the approximate Euler product representation
(12) $\frac{\Xi_{c}}{\Xi_{0}}(s)=\kappa_{0}\cdot\prod_{k=1}^{\infty}\exp\!\bigl(-c\,e^{-ik\theta}\,p_{k}^{-s}\bigr)\cdot M(s),$
where $\kappa_{0}=e^{cC_{1}}$ is an explicit nonzero constant (with $C_{1}=C_{1}(\theta)$ the non-oscillatory constant from (10)), and $M(s)$ is bounded analytic with $M(s)$ bounded for $\sigma>1/2$, $|t|\geq 2$. The Euler product $\prod\exp(-ce^{-ik\theta}p_{k}^{-s})=\sum a_{n}\,n^{-s}$ has $|a_{p_{k}}|=c$, $|a_{p_{k}^{m}}|=c^{m}/m!$, and $a_{mn}=a_{m}a_{n}$ for $\gcd(m,n)=1$.
Proof.
The spectral identity (8) gives $\log(\Xi_{c}/\Xi_{0})(s)=-cD(s)+R(s)$ with $R$ bounded. Write R(s)=R(2)+(R(s)R(2))R(s)=R(2)+(R(s)-R(2)); the constant $\kappa_{0}=e^{R(2)}$ is a nonzero number and $M(s)=e^{R(s)-R(2)}$ is bounded analytic. Then $\Xi_{c}/\Xi_{0}=\kappa_{0}\cdot\exp(-cD(s))\cdot M(s)$. Expanding the exponential of the prime-supported Dirichlet series gives the Euler product and multiplicative coefficients. ∎
Since both $\Xi_{c}$ and $\Xi_{0}$ satisfy $\Xi(s)=\Xi(1\!-\!s)$, the ratio inherits the functional equation $\Xi_{c}(s)/\Xi_{0}(s)=\Xi_{c}(1-s)/\Xi_{0}(1-s)$. The ratio is meromorphic: it has zeros at $\{1/2+i\gamma_{n}\}$ (from $\Xi_{c}$) and poles at $\{1/2+i\gamma_{n}^{(0)}\}$ (from $\Xi_{0}$), all on $\sigma=1/2$.
6. Almost-Periodicity
The Dirichlet series $D(s)$ itself has a clean almost-periodicity hierarchy.
Proposition 6.1.
  1. (i)
    For $\sigma>1$: $t\mapsto D(\sigma+it)$ is Bohr almost-periodic (uniformly AP), since $D$ is an absolutely convergent Dirichlet series.
  2. (ii)
    For $\sigma>\tfrac{1}{2}$: $t\mapsto D(\sigma+it)$ is Besicovitch ($B^{2}$) almost-periodic with $\|D(\sigma+i\,\cdot\,)\|_{B^{2}}^{2}=\sum_{k=1}^{\infty}p_{k}^{-2\sigma}<\infty$.
  3. (iii)
    The restriction $E(T)=\pi^{-1}\operatorname{Im}\,D(\tfrac{1}{2}+iT)$ is not $B^{2}$-AP: $\sum p_{k}^{-1}=\infty$ (Mertens).
Proof.
(i) Absolute convergence of $\sum p_{k}^{-\sigma}$ for $\sigma>1$ gives uniform convergence of the trigonometric series, hence Bohr-AP. (ii) The Fourier–Bohr coefficients of $D(\sigma+it)$ at exponents $\omega_{k}=\log p_{k}$ have squared moduli $p_{k}^{-2\sigma}$. Since $\sum p_{k}^{-2\sigma}$ converges for $\sigma>1/2$, the Riesz–Fischer theorem gives the result. (iii) Since $\sum p_{k}^{-1}=\infty$ (Mertens), the Fourier energy diverges at $\sigma=1/2$. ∎
The spectral identity transfers these properties to $\log(\Xi_{c}/\Xi_{0})$ up to the bounded error $R(s)$. The Fourier–Bohr coefficients of $\log(\Xi_{c}/\Xi_{0})$ at frequencies $\omega_{k}$ are $c\,e^{-ik\theta}\,p_{k}^{-\sigma}$ to leading order; since $R$ is bounded but need not itself be almost-periodic, the AP properties of $\log(\Xi_{c}/\Xi_{0})$ hold to leading order.
The AP hierarchy mirrors that of $\zeta$: Bohr-AP for $\sigma>1$ (Bohr [1]), $B^{2}$-AP for $\sigma>1/2$ (Bohr–Jessen [3]), with the convergent-to-divergent Fourier norm transition at $\sigma=1$.
7. Observations
We illustrate the construction numerically with $c=0.05$ and a finite truncation $K=80$. The computations use the shifted variant $\alpha_{k}=p_{k}+\theta$ in place of $p_{k}$, which ensures global monotonicity of $F_{K}(T)=N_{\mathrm{main}}(T)+cE_{K}(T)$ (see Remark 7.1). Since $\alpha_{k}\sim p_{k}$ as $k\to\infty$, the convergence/divergence properties at $\sigma=1$ are unchanged, and the spectral identity has the same structure with $\alpha_{k}$ replacing $p_{k}$.
Remark 7.1.
For the finite truncation $E_{K}$, global monotonicity of $F_{K}$ holds when $c<N_{\mathrm{main}}^{\prime}(T_{0})/(\sum_{k=1}^{K}\omega_{k}\alpha_{k}^{-1/2})$, which is satisfied for the parameters used. For the full infinite series, occasional crossings ($\gamma_{n+1}<\gamma_{n}$) may occur but do not affect the Hadamard product or the spectral identity (which depend on the linearization at $\gamma_{n}^{(0)}$, not on the ordering of $\gamma_{n}$).
7.1. First zeros of $Z_{c}$
The perturbed zeros $\gamma_{n}$ (defined by $F_{K}(\gamma_{n})=n$ for the truncation) and their shifts $\delta_{n}=\gamma_{n}-\gamma_{n}^{(0)}$:
$n$ $\gamma_{n}$ $\delta_{n}$ gap $E(\gamma_{n})$
1 17.9157 +0.068+0.068 0.011-0.011
2 23.2341 +0.062+0.062 5.318 0.013-0.013
3 27.7063 +0.035+0.035 4.472 0.008-0.008
4 31.6613 0.057-0.057 3.955 +0.015+0.015
5 35.4981 +0.030+0.030 3.837 0.008-0.008
10 51.7580 +0.024+0.024 3.061 0.008-0.008
20 78.6866 +0.014+0.014 2.522 0.006-0.006
50 144.5359 +0.005+0.005 2.038 0.003-0.003
100 236.8981 +0.045+0.045 1.808 0.026-0.026
Max shift: $|\delta_{n}|\leq 0.077$. Min gap: $1.67>0$. Gaps decrease as $\sim 2\pi/\log(\gamma_{n}/(2\pi))$, matching the Weyl law.
7.2. Zero statistics: theory and numerics
The microscopic statistics of $\Xi_{c}$’s zeros can be analyzed both theoretically and numerically. The key quantity is the ratio of perturbation to spacing:
(13) $\frac{|\delta_{n}|}{\text{spacing}}\approx c\,|E(\gamma_{n}^{(0)})|.$
Mean square of $E$ at the zeros. By Weyl equidistribution of the phases $\{\omega_{k}\gamma_{n}^{(0)}+\phi_{k}\}$ modulo $2\pi$ (for each finite truncation $K$, then taking $K\to\infty$):
(14) $\frac{1}{N}\sum_{n=1}^{N}E(\gamma_{n}^{(0)})^{2}\;\longrightarrow\;\frac{1}{2\pi^{2}}\sum_{k=1}^{\infty}p_{k}^{-1}\;=\;\frac{\log\log N+M}{2\pi^{2}}+o(1)$
(where $M$ is the Meissel–Mertens constant), so the RMS of $E(\gamma_{n}^{(0)})$ grows like $\sqrt{\log\log N/(2\pi^{2})}$. Consequently:
  1. (1)
    RMS of the counting error: $\bigl(\frac{1}{N}\sum(N_{c}(\gamma_{n})-n)^{2}\bigr)^{1/2}\sim c\sqrt{\log\log N/(2\pi^{2})}$, which matches the growth rate of $S(T)$ for $\zeta$ (cf. [17, Thm. 14.24]) but scaled by $c$.
  2. (2)
    Perturbation-to-spacing ratio: By (13), the RMS ratio is $c\sqrt{\log\log N/(2\pi^{2})}$, which exceeds $1$ only for $N>\exp\exp(2\pi^{2}/c^{2})$. For $c=0.05$ this threshold is $N>10^{10^{347}}$—far beyond any computation. Thus the zeros appear nearly crystalline (“super-rigid”) over any accessible range, but asymptotically the perturbations grow to many spacings, and the clock-like character is lost.
  3. (3)
    Normalized spacing variance: The deviation of the $n$-th normalized spacing from $1$ is $(\delta_{n+1}-\delta_{n})\cdot N_{\mathrm{main}}^{\prime}=c(E(\gamma_{n+1}^{(0)})-E(\gamma_{n}^{(0)}))$, which has variance $O(c^{2})$ for each $n$ (from the smoothness of $E$ on the scale of zero spacing). The normalized spacing std is therefore $O(c)$, much smaller than the GUE value $\approx 0.42$.
Numerical observations ($c=0.05$, $K=80$, 1000 zeros with $t\in[17.9,1420.1]$):
Statistic $\zeta$ (conditional on RH) $Z_{c}$ (observed)
RMS of $N-N_{\mathrm{main}}$ $\sqrt{\frac{\log\log T}{2\pi^{2}}}\to\infty$ $0.013$
Spacing std (normalized) $\approx 0.42$ (GUE [12]) $0.020$
Number variance $\Sigma^{2}(L)$ $\sim\tfrac{2}{\pi^{2}}\log L$ $\lesssim 0.25$
The observed values reflect the regime $c|E|\ll 1$ and are consistent with the theoretical $O(c)$ spacing deviation. The asymptotic growth $c\sqrt{\log\log N}$ is invisible at these heights. The construction captures the macroscopic transition (bounded/unbounded, AP hierarchy) but not the microscopic GUE statistics of $\zeta$, which arise from the Euler product and prime correlations that $Z_{c}$ deliberately lacks.
7.3. Value distribution
Figure 1 shows the complex-plane images of $\log(\Xi_{c}/\Xi_{0})(\sigma+it)$, $\log\zeta(\sigma+it)$, and $\log L(\sigma+it,\chi_{4})$ as $t$ ranges over $[20,300]$, for $\sigma=1.25,1.0,0.75$.
By the spectral identity and the $B^{2}$-AP structure of $D$, the mean square of $\log(\Xi_{c}/\Xi_{0})$ satisfies (to leading order in $c$): $\langle(\operatorname{Re}\log\frac{\Xi_{c}}{\Xi_{0}})^{2}\rangle\approx\langle(\operatorname{Im}\log\frac{\Xi_{c}}{\Xi_{0}})^{2}\rangle\approx\tfrac{c^{2}}{2}\sum p_{k}^{-2\sigma}$ (to leading order, from the $B^{2}$-norm of $D$). The value distribution is isotropic in the complex plane, as for $\log\zeta$ via Bohr–Jessen [3].
The quantitative difference is dramatic: at $\sigma=0.75$, the RMS of $\log(\Xi_{c}/\Xi_{0})$ is approximately $0.046$, while $\log\zeta$ has RMS $\approx 0.88$. The construction’s value distribution is confined to a bounded disk, whereas $|\log\zeta(\sigma+it)|$ is unbounded. This means $\Xi_{c}/\Xi_{0}$ cannot satisfy Voronin universality [18], and its moments do not exhibit the $(\log T)^{k^{2}}$ growth of $\zeta$ (cf. [11]). The value distribution is isotropic but concentrated—another manifestation of the missing arithmetic.
Figure 1. Complex-plane image of $\log f(\sigma+it)$ for $t\in[20,300]$. Left: $\log(\Xi_{c}/\Xi_{0})$; center: $\log\zeta$; right: $\log L(s,\chi_{4})$. Rows: $\sigma=1.25,1.0,0.75$. Color encodes $t$ (yellow==early, purple==late). Note the ${\sim}20{\times}$ difference in scale between the left column and the others.
Figure 2. Shape comparison: $\log(\Xi_{c}/\Xi_{0})$ rescaled to match the RMS of $\log\zeta$ (left) vs. $\log\zeta$ (right). The rescaled curve fills a banded region (reflecting the coherent trigonometric structure) rather than the isotropic spirals of $\log\zeta$ (reflecting multiplicative independence of primes).
7.4. Numerical evidence for the linearization condition
Lemma 3.2 establishes that $\sum_{n=1}^{\infty}E(\gamma_{n}^{(0)})^{2}/n^{2}<\infty$. We verify key aspects numerically with $K=80$ terms and $N=5000$ zero positions.
Mean square and Weyl equidistribution. The empirical mean square agrees with the Weyl prediction:
NN $\frac{1}{N}\sum_{n=1}^{N}E(\gamma_{n}^{(0)})^{2}$ $\frac{1}{2\pi^{2}}\sum_{k=1}^{K}p_{k}^{-1}$ ratio
100 0.0688 0.0654 1.053
500 0.0668 0.0654 1.022
1000 0.0649 0.0654 0.994
5000 0.0648 0.0654 0.991
The ratio stabilizes near $1$, consistent with joint equidistribution of the $K$ phases $\{\omega_{k}\gamma_{n}^{(0)}+\phi_{k}\}$ modulo $2\pi$.
Spectral identity remainder. The remainder $R(t)$ in the spectral identity $\operatorname{Re}\log(\Xi_{c}/\Xi_{0})(\sigma\!+\!it)=-c\sum p_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+R(t)$ has:
$\sigma$ $\operatorname{mean}(R)$ $\operatorname{std}(R)$ $\max|R|$
1.50 0.015-0.015 $0.021$ $0.081$
1.10 0.016-0.016 $0.030$ $0.122$
1.01 0.016-0.016 $0.034$ $0.136$
1.00 0.016-0.016 $0.034$ $0.138$
The remainder is $O(c)$ as claimed, and the linearization contribution ($\sim 10^{-5}$) is negligible compared to the constant and lower-limit terms.
8. Open Problems and Concluding Remarks
The function $\Xi_{c}$ provides an explicit entire function of order $1$ that provably satisfies an analogue of RH (all zeros on $\sigma=1/2$) together with the sharp bounded/unbounded transition at $\sigma=1$ characteristic of $\zeta$. The ratio $\Xi_{c}/\Xi_{0}$ is, to leading order, a Dirichlet series with generalized Euler product (Proposition 5.1), providing a concrete object whose zeros are provably on the critical line and whose analytic properties can be studied unconditionally.
The construction demonstrates that the $\sigma=1$ transition is driven entirely by the marginal convergence/divergence of the Dirichlet series $D(s)=\sum e^{-ik\theta}p_{k}^{-s}$ at $\sigma=1$, in direct analogy with $\log\zeta(s)=P(s)+O(1)$ [17, §3.2]. In the context of the Selberg class, this shows that an RH-like zero distribution is compatible with the $\sigma=1$ transition—the two properties reinforce rather than obstruct each other.
Several directions for further investigation suggest themselves.
GUE statistics. The zeros of $Z_{c}$ are super-rigid (nearly crystalline spacing, §7.2), far from the GUE statistics of $\zeta$. What additional structure on the perturbation would force GUE pair correlation onto the zeros of $\Xi_{c}$?
Pointwise bounds on $E$. The mean square (14) shows that $|E(\gamma_{n}^{(0)})|\sim\sqrt{\log\log n}$ in RMS, but the worst-case bound (2) gives only $O(\gamma_{n}^{1/2+\varepsilon})$. Numerically, $|E(\gamma_{n}^{(0)})|\leq 1.05$ for all computed nn7.4). Proving that $|E(\gamma_{n}^{(0)})|=O(\sqrt{\log\log n})$ pointwise (matching the RMS rate) would sharpen the linearization estimate and clarify when the perturbations first exceed the zero spacing.
Growth rate on $\sigma=1$. The Phragmén–Lindelöf argument gives $\sup_{t}|Z_{c}(1+it)|=+\infty$ but not the rate. For $\zeta$, $|\zeta(1+it)|=\Omega(\log\log t)$ [17, Thm. 8.5]. Does $|Z_{c}(1+it)|$ achieve the same rate?
An actual Dirichlet series. The ratio $\Xi_{c}/\Xi_{0}$ is approximately a Dirichlet series with Euler product (Proposition 5.1), but the bounded correction factor $M(s)$ prevents it from being one exactly. The base function $\Xi_{0}$ is not a Dirichlet series either (its Fourier exponents $\log\gamma_{n}^{(0)}$ are not of the form $\log n$). Can one unconditionally construct a Dirichlet series $\sum a_{n}\,n^{-s}$ (convergent for $\sigma>1/2$) with all zeros provably on $\sigma=1/2$ and the $\sigma=1$ transition? The Davenport–Heilbronn example [4] shows that a functional equation alone does not suffice; some form of multiplicative structure appears necessary.
Acknowledgments
This paper was developed through a collaboration between the author and Claude (Anthropic). The author posed the original question (whether one can construct an explicit entire function with the Riemann Hypothesis and the $\sigma=1$ transition), proposed the unit-disc analogy that motivated the Hadamard product approach, identified that a family result is insufficient (insisting on a single function), and formulated the equidistribution criterion for unboundedness. Claude proposed the Hadamard product construction, developed the spectral identity and its proof, found the dyadic large-sieve argument that resolved the linearization convergence, suggested the $\alpha_{k}=p_{k}$ switch that yields the Dirichlet series structure, and wrote the paper.